OpenStudy (anonymous):

3^(x-5)=9 x=2 x=-3 x=7 x=8

OpenStudy (anonymous):

If you agree that 3^2 = 9 then we can also right it this way 3^(x-5) = 3^2 that means that x-5 = 2, so x = 7

sam (.sam.):

Create equivalent expressions in the equation that all have equal bases. 3^(x-5)=3^(2) Since the bases are the same, then two expressions are only equal if the exponents are also equal. (x-5)=2 Since -5 does not contain the variable to solve for, move it to the other side of the equation x=5+2 Add 2 to 5 to get7 x=7